If Joy's age now is 3 times her age 14 years ago, how old is she now?
Answer by checkley75(3666)

3X=X+14
2X=14
X=14/2
X=7 IS HER AGE 14 YEARS AGO THEREFORE SHE IS NOW 7+14=21 YEARS OLD.